This one is driving me up a wall?


The difference of two numbers is 33. The second is 7 less than 3 times the first. What are the two numbers?

Answer by Nate(3500)

The first number we will denote by: x
The second is 7 less than 3 times the first.
The second: 3x - 7
The difference of two numbers is 33
x - (3x - 7) = 33
or
(3x - 7) - x = 33
Usually depending on with number is higher ...
x - 3x + 7 = 33
or
3x - 7 - x = 33
~~
-2x = 26
or
2x = 40
~~
x = -13
or
x = 20
~~
x = -13
second number: 3(-13) - 7 = -39 - 7 = -46
or
x = 20
second number: 3(20) - 7 = 60 - 7 = 53